OpenStudy (anonymous):

find the rule for: 2, 7, 12, 17, 22, 27, 32

OpenStudy (anonymous):

you add 5 each time...

OpenStudy (anonymous):

well i got that, ha, but i have to write an equation for n.

OpenStudy (raden):

the difference of 2 terms is 5, so add 5 to get next terms

OpenStudy (anonymous):

oh it would be 5(n-1)+2 n being what number it is in the sequence

OpenStudy (anonymous):

thank you so much!

OpenStudy (raden):

that should like this, @Zorark an = a1 + (n-1)d here a1 = 2 and d = 5 so, an = a1 + (n-1)d = 2 + (n-1)5 = 2 + 5n - 5 = 5n - 3

OpenStudy (raden):

but, the last answer is same :)

OpenStudy (anonymous):

@RadEn thanks, can you please help me with my problem?

OpenStudy (raden):

ha, what is it ?

OpenStudy (anonymous):

What is the speed, in miles per hour of a point on the equator due to the rotation of earth? Assume that the earth is a sphere of radius 4000 miles, no ones answered it in an hour.. :(

OpenStudy (anonymous):

@RadEn